please help me to solve the following problem. m
Brian is 4 years older than Bob while Bob is 11 years older than Mark. If their combined age is 86 years, how old is Brian?

Answer by ankor@dixie-net.com(22740)

let x = Brian's age
let y = Bob's age
let z = Marks age
:
Write and equation for each phrase
:
Brian is 4 years older than Bob while
x = y + 4
Bob is 11 years older than Mark.
y = z + 11
or
z = y - 11
If their combined age is 86 years,
x + y + z = 86
Replace x and z using the above equations
(y+4) + y + (y-11) = 86
3y - 7 = 86
3y = 86 + 7
y = 93/3
y = 31
:
how old is Brian?
x = 31 + 4
x = 35 yrs is Brian
:
:
You can confirm this for yourself, find z and see if they add up to 86